OpenStudy (anonymous):

q 3/2 type the expression as a radical

OpenStudy (danjs):

\[q^{1/2} = \sqrt{q}\]

OpenStudy (danjs):

\[q^{3/2} = q^{1/2}*q^{2/2}\]